# thanks: Corresponding author

### v1

@@ III.1 Radio-load quasar population
×$\times\int\mathrm{d}M\frac{\mathrm{d}n(z,M)}{\mathrm{d}M}\int_{R_{0}(M,F_{\rm th})}\mathrm{d}RN(R),$
@@
where $R$ represents radio-loudness. $\bar{R}$ = 2.67 and $\sigma_{0}$ = 0.55 represent the best fitting parameters obtained from Ref. Niu et al. (2024).
To calculate the survey area of SKA-LOW, we apply the following formula,
@@
where $\eta$ represents the fraction of the initial survey area that remains usable for observation after excluding the Milky Way, for which we adopt a value of 0.7. Considering a declination range from $34^{\circ}$ to $-86^{\circ}$, the total survey area is approximately 10,000 deg2.
@@
where $\delta\nu$ is the channel width, $\delta t$ is the integration time, $A_{\rm eff}$ is the effective collecting area of the telescope and $T_{\rm sys}$ is the system temperature. In our analysis, we use $\delta\nu$ = 1 kHz and $\delta t$ = 100 h. We adopt $A_{\rm eff}/T_{\rm sys}=600\mathrm{~{}m}^{2}\mathrm{~{}K}^{-1}$ for the SKA-LOW Braun et al. (2019). We consider SNR greater than 5 as the criterion for detecting quasars. Assuming an optical depth $\tau$ = 0.1, to meet this criterion, we need to satisfy the condition $F_{\rm th}\times(1-e^{-\tau})>5\times\delta S^{\rm N}$.
We present the $z-S_{150}$ distribution of radio-loud quasars in Fig. 3. Assuming an integration time of $\delta t$ = 100 h and a channel width of $\delta\nu$ = 1 kHz, under this configuration, quasars with a flux $S_{150}\gtrsim$ 10 mJy at low redshifts between 6 and 8 can all be observed, and there are also a few quasars with a flux $S_{150}\gtrsim$ 100 mJy. As the redshift increases, only a small number of quasars with a flux above 10 mJy can be detected. Particularly, it becomes very difficult to find suitable background sources for observation at redshifts above 10. For these fainter quasars, observations can be improved by increasing the integration time or broadening the bandwidth. However, broadening the bandwidth reduces the number of measurements, which in $k$-space means fewer $k$ modes, leading to increased sample variance and thus greater uncertainty in the measurements. From these, we select 10 radio-loud quasars with significant flux densities at different redshifts for further analysis.
@@ III.2 Parameter estimation
$P^{N}=\frac{1}{\sqrt{N_{s}}}\left(\frac{\lambda_{z}^{2}T_{\mathrm{sys}}}{A_{\rm{eff}}\Omega}\right)^{2}\left(\frac{\Delta r_{z}}{2\Delta\nu_{z}\delta t_{0.5}}\right),$ (24)
where $N_{s}$ is the number of neutral segments measured, $\Delta v_{z}$ is the total observing bandwidth corresponding to $\Delta r_{z}$, $\delta t_{0.5}$ is half of the integration time. In actual observations of the 21-cm forest signal, we can divide the total observation time into two parts and perform cross-correlation on the signals observed in these two segments. This approach helps to reduce the thermal noise. Similarly, in our work, we employed this method to estimate the thermal noise. We utilize $P(k_{\|},z)$ to estimate the sample variance, and the total measurement error of the 1D power spectrum $\sigma_{P}$ is
$\sigma_{P}=\frac{1}{\sqrt{N_{\rm m}}}\left[P^{N}+P(k_{\|},z)\right],$ (25)
@@
Due to current observations only placing a lower limit on $m_{\rm W}$, we employ Bayesian analysis to deduce its reciprocal during the inference process. We employed Bilby Ashton et al. (2019) for Bayesian analysis to estimate the parameters $f_{\rm X}$ and $m_{\rm W}^{-1}$. For the observed 1D power spectrum $\tilde{P}$, it can be expressed as
@@
where $\sigma_{P_{i,k}}$ is the measurement error. The prior distributions for the parameters $f_{\rm X}$ and $m_{\rm W}^{-1}$ are uniform distributions over the range $[0.01,1]$ and $[1/9,1/3]~{}\rm{keV}^{-1}$. We used the Markov Chain Monte Carlo Ashton and Talbot (2021) sampling algorithm to draw samples from the posterior distribution, estimating the parameters $f_{\rm X}$ and $m_{\rm W}^{-1}$ and their precision.
@@ IV Results and discussion
Figure 4: Comparison of the 21-cm forest halo model presented in this paper (solid curves) with the simulation results (dashed curves) at $z$ = 9, assuming a flux density of $S_{150}$ = 10 mJy. The green, yellow and blue curves correspond to $f_{\rm X}$ = 0.01, 0.1 and 1, respectively. The left and right panels represent CDM model and WDM model with $m_{\rm W}=6$ keV, respectively.
@@
While the DM model and heating effects suppress the 21-cm forest signal, their impacts on different scales vary significantly. These differences can be revealed through the 1D power spectrum. To elucidate these scale-dependent effects, we developed an analytical model for the 1D power spectrum of the 21-cm forest using the halo model method and conducted detailed calculations of the 1D power spectra under various DM models and heating levels. We consider a distance of 2 comoving Mpc along the line of sight, corresponding to a bandwidth of $\Delta v_{z}=0.11~{}\mathrm{MHz}$ and $k_{\rm min}$ = 3.14 Mpc-1 at $z=10$. Here, we only consider modeling on small scales below 2 Mpc for comparison with the simulation result. Additionally, modeling on larger scales would involve considering the ionized fraction, which would add complexity to the model.
@@
For a direct comparison between our analytical model and the simulated 1D power spectra from previous studies, we conducted small-scale simulations on grids of 2 Mpc in length Shao et al. (2023). Each (2 Mpc)3 grid is populated with halos based on the conditional halo mass function for the CDM or WDM Cooray and Sheth (2002); Zentner (2007). The initial density is the mean density of the universe. The density in each voxel is determined by the NFW profile or the infall model profile. The gas temperature within each voxel is determined by its location relative to halos and the thermal history of the universe. The large-scale reionization history only determines the probability of encountering a neutral patch of the IGM of a certain length along the line of sight, and we focus only on neutral regions on small scales. Each grid is divided into 5003 voxels, and we average over 5002 lines along the line of sight direction to obtain the final 1D power spectrum for comparison with the analytical model, thereby mitigating the impact of sample variance.
@@
In Fig. 4, we present a comparison between the 21-cm forest halo model developed in this work and the simulation results obtained using methods from Ref. Shao et al. (2023). We display the results for both the CDM model (left panel) and the WDM model (right panel) at different heating levels, with fixed redshift $z$ = 9 and $S_{150}$ = 10 mJy. Our analytical model shows good agreement with simulation results across various DM models and different levels of X-ray heating. It is noteworthy that, at such small scales, the one-halo term dominates the 1D power spectrum. Therefore, we only present the results for the one-halo term.
@@
Figure 5: Comparison between the 21-cm forest halo model in this paper (solid curves) and the simulation results (dashed curves). The blue, yellow and green curves in left panel correspond to $z$ = 6, 8 and 10, respectively, with $S_{150}$ = 10 mJy. The green, yellow and blue curves in right panel correspond to $S_{150}$ = 100 mJy, 10 mJy and 1 mJy, respectively, at $z$ = 9.
@@
We fix $m_{\rm W}$ = 6 keV and $f_{\rm X}$ = 0.1 as the fiducial model, and calculate the performance of the 21-cm forest halo model at different redshifts and with different background source flux densities in Fig. 5. The left panel shows the results for redshifts $z$ = 6, 8 and 10, assuming a uniform flux density $S_{150}$ = 10 mJy for consistency. Note that in practice, at low redshifts, many quasars can have a flux density exceeding 10 mJy, whereas at high redshifts, such quasars are relatively few. The right panel fixes the redshift $z$ = 9 and calculates the results for different background source flux densities $S_{150}$ = 1 mJy, 10 mJy and 100 mJy. This indicates that our model is applicable across a wide range of parameter space.
@@
To accurately describe the heating history of the universe and constrain the heating level $f_{\rm X}$, we selected 10 quasars, each with 10 segments, at different redshifts. The SNRs of the 1D power spectra of the 21-cm forest for these quasars are shown in Fig. 6. Considering that thermal noise varies with redshift, we present the SNR, $P(k)/P^{N}$. At higher redshifts, the temperature of the IGM is lower, which means that quasars can achieve sufficient SNR even with moderate flux densities. In contrast, at lower redshifts, quasars need to be brighter to attain the same level of SNR. Quasars at redshifts between 7 and 8 are particularly well-suited for observations of the 1D power spectrum of the 21-cm forest.
At the end of the EoR, the neutral hydrogen fraction of the universe is already quite low, and most of the universe at $z\sim 6$ has been ionized. However, since the lengths of the segments we are considering are very short, only about 2 Mpc, it is still possible to have up to ten segments of neutral regions. Additionally, for quasars at higher redshifts, there may be many more neutral segments. To simplify, we assume that each quasar has 10 segments, totaling 100 segments. We take the redshift of the quasars as the redshift of the 21-cm forest signal and assume that the quasars do not produce a heating effect.
@@
We applied the Bayesian methods to constrain the parameters $m_{\rm W}$ and $f_{\rm X}$, which are presented in Fig. 7. Based on the prospective observations with SKA-LOW, we assumed 100 h of observation for each quasar, accumulating to a total of 1000 h of observational time. We adopted a fiducial model with $m_{\rm W}$ = 6 keV and $f_{\rm X}$ = 0.1, and calculated the absolute errors for the parameters to be $\sigma_{m_{\rm W}}$ = 1.3 keV and $\sigma_{f_{\rm X}}$ = 0.02. By combining 21-cm forest data from various redshifts, we can effectively constrain the mass of WDM particles and trace the evolution of the cosmic thermal history, thereby revealing the formation of the first galaxies.
@@
Observing the 21-cm forest is contingent upon high-redshift, radio-bright background sources, with radio-loud quasars and radio afterglows of gamma-ray bursts emerging as the most viable options. As some of the brightest objects in the universe, quasars’ radio emissions can serve as the background for the 21-cm forest. In particular, radio-loud quasars provide sufficient intensity for observing the 21-cm signal. Using the quasar luminosity function, it is possible to predict the number of quasars, and in the SKA era, it is expected that a sufficient number of quasars will be detected to serve as background sources for the 21-cm forest Banados et al. (2021); Ighina et al. (2021); Haiman et al. (2004); Gloudemans et al. (2022). Some models have suggested that although the abundance of radio-loud quasars at high redshifts is low, it is still adequate to support 21-cm forest studies Gloudemans et al. (2021); Niu et al. (2024).
V Conclution

### v2

@@ III.1 Radio-loud quasar population
$\times\int\mathrm{d}M\frac{\mathrm{d}n(z,M)}{\mathrm{d}M}\int_{R_{0}(M,F_{\rm th})}\mathrm{d}RN(R),$
@@
where $R$ represents radio loudness. $\bar{R}=2.67$ and $\sigma_{0}=0.55$ represent the best fitting parameters obtained from Ref. Niu et al. (2024).
To calculate the survey area of SKA-LOW, we apply the following formula:
@@
where $\eta$ represents the fraction of the initial survey area that remains usable for observation after excluding the Milky Way, for which we adopt a value of 0.7. Considering a declination range from $34^{\circ}$ to $86^{\circ}$, the total survey area is approximately 10,000 deg2.
@@
where $\delta\nu$ is the channel width, $\delta t$ is the integration time, $A_{\rm eff}$ is the effective collecting area of the telescope and $T_{\rm sys}$ is the system temperature. In our analysis, we use $\delta\nu=1{\rm~kHz}$ and $\delta t=100{\rm~h}$. We adopt $A_{\rm eff}/T_{\rm sys}=600\mathrm{~m}^{2}\mathrm{~K}^{-1}$ for the SKA-LOW Acedo et al. (2020); Braun et al. (2019). We consider SNR greater than 5 as the criterion for detecting radio-loud quasars. Assuming an optical depth $\tau=0.1$, to meet this criterion, we need to satisfy the condition $F_{\rm th}\times(1-e^{-\tau})>5\times\delta S^{\rm N}$.
We present the $z-S_{150}$ distribution of radio-loud quasars in Fig. 3. Assuming an integration time of $\delta t=100{\rm~h}$ and a channel width of $\delta\nu=1{\rm~kHz}$, under this configuration, quasars with a flux $S_{150}\gtrsim 10{\rm~mJy}$ at low redshifts between 6 and 8 can all be observed, and there are also a few quasars with a flux $S_{150}\gtrsim 100{\rm~mJy}$. As the redshift increases, only a small number of quasars with a flux above $10{\rm~mJy}$ can be detected. Particularly, it becomes very difficult to find suitable background sources for observation at redshifts above 10. For these fainter quasars, observations can be improved by increasing the integration time or decreasing the frequency resolution. However, lower frequency resolution reduces the number of measurements, which in $k$ space means fewer $k$ modes, leading to increased sample variance and thus greater uncertainty in the measurements. From these, we select 10 radio-loud quasars with significant flux densities at different redshifts for further analysis.
@@ III.2 Parameter estimation
$P^{\rm N}=\left(\frac{\lambda_{z}^{2}T_{\mathrm{sys}}}{A_{\rm{eff}}\Omega}\right)^{2}\left(\frac{\Delta r_{z}}{2\Delta\nu_{z}\delta t_{0.5}}\right),$ (24)
where $\Delta\nu_{z}$ is the total observing bandwidth corresponding to $\Delta r_{z}$, $\delta t_{0.5}$ is half of the integration time. Here, $\Delta r_{z}$ denotes the comoving length of the observed segment along the line of sight. If $N_{s}$ neutral segments can be observed, the thermal noise will be reduced by a factor of $\sqrt{N_{s}}$. Here, $N_{s}$ represents the total number of usable neutral segments along the sightlines of all observed quasars, which depends on the number of background sources and the number of neutral regions that can be segmented in the spectrum of each background source. In actual observations of the 21-cm forest signal, we can divide the total observation time into two parts and perform cross-correlation on the signals observed in these two segments. This approach helps to reduce the thermal noise. Similarly, in our work, we employed this method to estimate the thermal noise. We utilize $P(k_{\|},z)$ to estimate the sample variance, and the total measurement error of the 1D power spectrum $\sigma_{P}$ is
$\sigma_{P}=\frac{1}{\sqrt{N_{\rm m}}}\left[P^{\rm N}+P(k_{\|},z)\right],$ (25)
@@
Because of current observations only placing a lower limit on $m_{\rm W}$, we employ Bayesian analysis to deduce its reciprocal during the inference process. We employed Bilby Ashton et al. (2019) for Bayesian analysis to estimate the parameters $f_{\rm X}$ and $m_{\rm W}^{-1}$. For the observed 1D power spectrum $\tilde{P}$, it can be expressed as
@@
where $\sigma_{P_{i,k}}$ is the measurement error. The prior distributions for the parameters $f_{\rm X}$ and $m_{\rm W}^{-1}$ are uniform distributions over the range $[0.01,1]$ and $1/9,1/3]~\rm{keV}^{-1}$. We used the Markov chain Monte Carlo Ashton and Talbot (2021) sampling algorithm to draw samples from the posterior distribution, estimating the parameters $f_{\rm X}$ and $m_{\rm W}^{-1}$ and their precision.
@@ IV Results and discussion
Figure 4: Comparison of the 21-cm forest halo model presented in this paper (solid curves) with the simulation results (dashed curves) at $z=9$, assuming a flux density of $S_{150}=10{\rm~mJy}$. The green, yellow and blue curves correspond to $f_{\rm X}$ = 0.01, 0.1 and 1, respectively. The left and right panels represent CDM model and WDM model with $m_{\rm W}=6{\rm~keV}$, respectively.
@@
While the DM model and heating effects suppress the 21-cm forest signal, their impacts on different scales vary significantly. These differences can be revealed through the 1D power spectrum. To elucidate these scale-dependent effects, we developed an analytical model for the 1D power spectrum of the 21-cm forest using the halo model method and conducted detailed calculations of the 1D power spectra under various DM models and heating levels. We consider a distance of 2 comoving Mpc along the line of sight, corresponding to a bandwidth of $\Delta\nu_{z}=0.11~\mathrm{MHz}$ and $k_{\rm min}=3.14~{\rm Mpc}^{-1}$ at $z=10$. Here, we only consider modeling on small scales below $2{\rm~Mpc}$ for comparison with the simulation result. Additionally, modeling larger scales that correspond to a broader bandwidth requires consideration of the ionized fraction, which adds complexity to the model. For some radio-loud quasars at relatively low redshifts, especially those near redshift 6 during the late stages of reionization, the ionization fraction of the Universe has already increased significantly. As a result, neutral regions are separated by large ionized regions. Against this backdrop, it is still feasible to extract short neutral segments from quasar spectra for the analysis of the 21-cm forest. However, extracting longer lines to probe larger-scale structures becomes very challenging due to the presence of ionized regions. The ionization field not only affects the length and distribution characteristics of the neutral segments but also impacts the 1D power spectrum of the 21-cm forest. Therefore, to extend our analytical model to larger scales, it is necessary in future studies to incorporate the effects of ionization into the modeling framework. On the other hand, a larger bandwidth also increases the number of accessible $k$ modes, which can help reduce sample variance in statistical measurements.
@@
For a direct comparison between our analytical model and the simulated 1D power spectra from previous studies, we conducted small-scale simulations on grids of $2{\rm~Mpc}$ in length Shao et al. (2023). Each $(2{\rm~Mpc})^{3}$ grid is populated with halos based on the conditional halo mass function for the CDM or WDM Cooray and Sheth (2002); Zentner (2007). The initial density is the mean density of the Universe. Specifically, within halos, the gas temperature is set to the virial temperature, while in the surrounding IGM, it is determined by adiabatic cooling or x-ray heating. The heating rate is regulated by an efficiency parameter $f_{\rm X}$, which controls the strength of x-ray emission from early x-ray sources. The density in each voxel is determined by the NFW profile or the infall model profile. The gas temperature within each voxel is determined by its location relative to halos and the thermal history of the Universe. The large-scale reionization history only determines the probability of encountering a neutral patch of the IGM of a certain length along the line of sight, and we focus only on neutral regions on small scales. We assume that gas within and around low-mass halos is in collisional ionization equilibrium. Since the ionization fraction in such regions remains very low, we neglect ionization effects and consider only fully neutral patches. After generating the density and temperature fields in the simulation, we further assume different background source flux densities to compute the 21-cm forest brightness temperature and the corresponding 1D power spectrum. Each grid is divided into $500^{3}$ voxels, corresponding to a voxel size of approximately 4 kpc, and we average over $500^{2}$ lines along the line of sight direction to obtain the final 1D power spectrum for comparison with the analytical model, thereby mitigating the impact of sample variance. It is worth noting that the multiscale simulations presented in Ref. Shao et al. (2023) combine large-scale reionization and density fields from 21cmFAST with small-scale modeling of halo properties. In contrast, our analytical model focuses solely on small-scale structures. Therefore, for consistency, we compare our results only with the small-scale parts (i.e., scales below 2 Mpc) of the simulations in Ref. Shao et al. (2023).
@@
In Fig. 4, we present a comparison between the 21-cm forest halo model developed in this work and the simulation results obtained using methods from Ref. Shao et al. (2023). We display the results for both the CDM model (left panel) and the WDM model (right panel) at different heating levels, with fixed redshift $z=9$ and $S_{150}=10{\rm~mJy}$. Our analytical model shows good agreement with simulation results across various DM models and different levels of x-ray heating. It is noteworthy that, at such small scales, the one-halo term dominates the 1D power spectrum. Therefore, we only present the results for the one-halo term.
@@
Figure 5: Comparison between the 21-cm forest halo model in this paper (solid curves) and the simulation results (dashed curves). The blue, yellow, and green curves in the left panel correspond to $z$ = 6, 8, and 10, respectively, with $S_{150}=10{\rm~mJy}$. The blue, yellow, and green curves in the right panel correspond to $S_{150}=1$, $10$, and $100{\rm~mJy}$, respectively, at $z=9$.
@@
We fix $m_{\rm W}=6{\rm~keV}$ and $f_{\rm X}=0.1$ as the fiducial model, and calculate the performance of the 21-cm forest halo model at different redshifts and with different background source flux densities in Fig. 5. The left panel shows the results for redshifts $z$ = 6, 8, and 10, assuming a uniform flux density $S_{150}=10{\rm~mJy}$ for consistency. Note that in practice, at low redshifts, many quasars can have a flux density exceeding $10{\rm~mJy}$, whereas at high redshifts, such quasars are relatively few. The right panel fixes the redshift $z=9$ and calculates the results for different background source flux densities $S_{150}=1$, $10$, and $100{\rm~mJy}$, respectively. This indicates that our model is applicable across a wide range of parameter space.
@@
To accurately describe the heating history of the Universe and constrain the heating level $f_{\rm X}$, we aimed to identify the most promising radio-loud quasars for 21-cm forest observations. We first computed the total signal-to-noise ratio, $\sum P(k)/P^{\rm N}$, of the 1D power spectrum of the 21-cm forest for all radio-loud quasars detectable by SKA. Based on this, we selected the 10 sources with the highest total SNRs for our analysis. This selection reflects both an idealized and realistic scenario: under limited observational resources, prioritizing brighter quasars with higher SNR allows for more effective constraints on key physical parameters, such as the mass of warm dark matter particles and the level of cosmic heating. Such an approach is not only theoretically motivated but also aligns with realistic time-allocation considerations in actual observations. For each of these 10 quasars, we assume the presence of 10 neutral segments along the line of sight, each with a comoving length of 2 Mpc. The resulting SNRs of the 1D power spectra are shown in Fig. 6. Considering that thermal noise varies with redshift, we present the SNR in the form $P(k)/P^{\rm N}$. At higher redshifts, the temperature of the IGM is lower, which means that quasars can achieve sufficient SNR even with moderate flux densities. In contrast, at lower redshifts, quasars need to be brighter to attain the same level of SNR. Quasars at redshifts between 7 and 8 are particularly well suited for observations of the 1D power spectrum of the 21-cm forest.
At the end of the EoR, the neutral hydrogen fraction of the Universe is already quite low, and most of the Universe at $z\sim 6$ has been ionized. However, numerous observations suggest that the end of reionization occurred later than anticipated, possibly close to $z\sim 5.3$ Kulkarni et al. (2019); Nasir and D’Aloisio (2020); Qin et al. (2021). Therefore, it is also possible to detect a significant amount of 21-cm forest signals during the late stages of reionization Šoltinský et al. (2021). Furthermore, since the lengths of the segments we are considering are very short, only about $2{\rm~Mpc}$, it is still possible to have up to ten segments of neutral regions. Additionally, for radio-loud quasars at higher redshifts, there may be many more neutral segments. To simplify, we assume that each quasar has 10 segments, totaling 100 segments. We take the redshift of the quasars as the redshift of the 21-cm forest signal and assume that the quasars do not produce a heating effect.
@@
We applied the Bayesian methods to constrain the parameters $m_{\rm W}$ and $f_{\rm X}$, which are presented in Fig. 7. Based on the prospective observations with SKA-LOW, we assumed 100 h of observation for each quasar, accumulating to a total of 1000 h of observational time. We adopted a fiducial model with $m_{\rm W}=6{\rm~keV}$ and $f_{\rm X}=0.1$, and calculated the absolute errors for the parameters to be $\sigma_{m_{\rm W}}=1.3{\rm~keV}$ and $\sigma_{f_{\rm X}}=0.02$. By combining 21-cm forest data from various redshifts, we can effectively constrain the mass of WDM particles and trace the evolution of the cosmic thermal history, thereby revealing the formation of the first galaxies.
@@
Observing the 21-cm forest is contingent upon high-redshift, radio-bright background sources, with radio-loud quasars and radio afterglows of gamma-ray bursts emerging as the most viable options. As some of the brightest objects in the Universe, quasars’ radio emissions can serve as the background for the 21-cm forest. In particular, radio-loud quasars provide sufficient intensity for observing the 21-cm signal. Using the quasar luminosity function, it is possible to predict the number of quasars, and in the SKA era, it is expected that a sufficient number of radio-loud quasars will be detected to serve as background sources for the 21-cm forest Banados et al. (2021); Ighina et al. (2021); Haiman et al. (2004); Gloudemans et al. (2022). Some models have suggested that, although the abundance of radio-loud quasars at high redshifts is low, it is still adequate to support 21-cm forest studies Gloudemans et al. (2021); Niu et al. (2024).
It is commonly believed that galaxies can form only in dark matter halos with masses above $10^{8}M_{\odot}$. However, some studies have suggested that Pop III galaxies might also form in low-mass halos with masses below $10^{8}M_{\odot}$ via molecular cooling. The ultraviolet and x-ray radiation produced by these galaxies can heat and ionize the IGM, thereby affecting the statistical properties of the 21-cm signal Pochinda et al. (2024); Lazare et al. (2024). Nevertheless, the halo model adopted in this paper is primarily designed for the statistical modeling of the 21-cm forest absorption signal structure and does not yet account for the potential formation of galaxies in low-mass halos and their radiative feedback. The formation of molecular-cooling galaxies in low-mass halos and their radiative processes, especially their effects on the 21-cm forest signal and the 1D power spectrum in particular, are important topics that require further in-depth investigation in the future.
V Conclusion